test the null hypothesis H0

H0 : In college, the average GPA of men is equal to the average GPA of women.
H1 : In college, the average GPA of men is different from the average GPA of women.

A sample of 10 men's GPA in college has sample mean 2.9, and a sample of 10 women's GPA has
sample mean 3.1. We also know the GPAs of men and women have the same standard deviation
0.2. Calculate the p value.

Please show your work so I can follow along, thanks!
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: μ1 = μ2
Alternative Hypothesis, Ha: μ1 ≠ μ2

Pooled Variance
sp = sqrt(s1^2/n1 + s2^2/n2)
sp = sqrt(0.04/10 + 0.04/10)
sp = 0.0894

Test statistic,
z = (x1bar - x2bar)/sp
z = (2.9 - 3.1)/0.0894
z = -2.24

P-value Approach
P-value = 0.0251
As P-value < 0.1, reject null hypothesis.


The two means are not equal